OpenStudy (anonymous):

(a4/3 over a1/2) 2/5 to the side

OpenStudy (anonymous):

\[{a{4\over3} \over a{1\over2}}={2\over5} \] ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

It says to perform the operations and simplify. Leave answers exponential form.

OpenStudy (lgbasallote):

i think there's no =

OpenStudy (anonymous):

4a/3/a/2=2/5, 3a/8a=2/5,

OpenStudy (lgbasallote):

\[\huge \left(\frac{a^{\frac 43}}{a^\frac 12} \right) \frac 25\]

OpenStudy (anonymous):

ahh...

OpenStudy (anonymous):

yep that's it Igbasallote

OpenStudy (anonymous):

(3a/8a)^2/5,

OpenStudy (anonymous):

how you got that

OpenStudy (anonymous):

nvm it's wrong

OpenStudy (anonymous):

first, you can multiply 2/5 to both 4/3 and 1/2

OpenStudy (anonymous):

a^(8/15)/a^(2/10)

OpenStudy (anonymous):

10.12.2

OpenStudy (anonymous):

this the last one

OpenStudy (anonymous):

this was some straight pellet in the book smh